• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Abdurahim

    Log entry time 20:16:30 on February 9,2011

    Entry number 346241

    keyword=Target gui & charts

    L1 & L2 temperatures have been stablized after Dave turned on heaters and possibly adjusted the PID loop. But the read backs don't seem to be working very well. Some of them are unbelievably smooth. Take a look at charts.

    FIGURE 1

    FIGURE 2